find the equation of each line (by its coressponding number)
ارو

you can use slope formula format for each
Answer #1

We use the slope formula

( y - y1 ) / (y2 - y1) = ( x - x1 ) / ( x2 - x1 )

We get the following equations :

Solution : Eqr of (4,6) x-2 Eqr of (7,10) line 1: (2, 10) to Y-10 6-10 4-2 y-10 = – 2(x-2) (2x + y = 14 line 2 : (5, 6) to yEar of line (-7,7) 93 (-4,-1) to Y+1 = x+4 7+ 1 -7+4 3y + 3 = - 8x - 32 8x+3y = – 35 Eqr of line 10: (-4,-1) to (41) yti Y+1Eqr of line is : (5.5) to (14,8) 3(4-5) = (3y - x = x-5 10 Ear of line 16 : (14,8) to (16,7) y - 8 7 - x - 14 16- 14 8 Ear ofEqr of (12,-2) to (11,-6) line 20 : y + 2 - 6+2 x - 12 - 12 972 = x-12 - 4 y+2 1 4x + y = – 4x + 48 = 46 Eqr of Ear of line 2